# Upper and Lower Bounds on the Smoothed Complexity of the Simplex Method

### v2

@@ 5.1 Notations
Since we assume that $a_{1},\dots,a_{n}$ each have a continuous probability density function, $\operatorname{conv}(a_{1},\dots,a_{n})$ and $W$ satisfy the non-degenerate conditions (see 23) almost surely. In this case, each edge of the polygon $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ is given by a $F_{I}\cap W=\operatorname{conv}(a_{i}:i\in I)$ for which $I\in\binom{[n]}{d}$ and $E_{I}$ holds (where $F_{I}$ and $E_{I}$ are defined in Definition 25). In addition, each vertex of the polygon $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ is given by the intersection between $W$ and $(d-2)$-dimensional ridges of $\operatorname{conv}(a_{1},\ldots,a_{n})$, which are convex hulls of $(d-1)$ vertices of $\operatorname{conv}(a_{1},\ldots,a_{n})$. We define the following notations of ridge and the corresponding vertex:
Definition 34 (Ridge and vertex event).
@@
Remark 35.
Any vertex $v$ of $\operatorname{conv}(a_{1},\dots,a_{n})$ can be written as $v=R_{J}\cap W$ for some $J\subseteq[n]$ for which $A_{J}$ holds. Assuming non-degeneracy, each $J$ for which $A_{J}$ holds satisfies $|J|=d-1$ and the relation between vertices and index sets $J\in\binom{[n]}{d-1}$ with $A_{J}$ is a one-to-one correspondence.
@@ 5.2 Deterministic Conditions for a Good Edge-to-Vertex Separator
In this subsection, we present a series of sufficient conditions that an edge on the polygon $\operatorname{conv}(a_{1},\cdots,a_{n})\cap W$ maintains a significant distance to its next vertex in the clockwise order.
Lemma 36.
@@
    (Inner radius of $R$) $\operatorname{dist}(F\cap W,\partial R)\geq r$.
Then the outer unit normal vector $\bar{\theta}\in W$ to the edge $F\cap W$ satisfies
$\bar{\theta}^{\top}p-\bar{\theta}^{\top}p^{\prime}\geq\delta r/3$
for any $p\in F\cap W$, here $p^{\prime}\in Q\cap W$ is the next vertex after $F\cap W$ in clockwise order.
We remark that Lemma 36 only provides a deterministic characterization of $(a_{1},\cdots,a_{n})$’s that result in large edge-to-vertex distance. However, to show that such edge-to-vertex distance exist with good probability, we need to further explore the characteristics of $(a_{1},\cdots,a_{n})$’s distributions.
Proof.
Write $R^{\prime}$ for the ridge of $Q$ such that $\{p^{\prime}\}=R^{\prime}\cap W$. Since $p^{\prime}\in Q\cap W$ is adjacent to vertex $p$ and the edge $F\cap W$, by 24 we may relabel the $a_{i}$ such that $R^{\prime}=\operatorname{conv}(a_{1},\dots,a_{d-1})$, $R=\operatorname{conv}(a_{2},\dots,a_{d}),$ and $F=\operatorname{conv}(a_{2},\dots,a_{d+1})$ without loss of generality. Let $\theta\in\mathbb{S}^{d-1}$ denote the outward unit normal to $F$. This normal vector satisfies
Let $s\in\mathbb{S}^{d-1}$ be the unit vector indicating the direction of the (one-dimensional) line $F\cap W$. This vector is unique up to sign. Also, let $\bar{\theta}=\pi_{W}(\theta)/\|\pi_{W}(\theta)\|$ be the outward unit normal to $F\cap W$ in the two-dimensional plane $W$. Notice that $\bar{\theta}$ and $s$ form an orthonormal basis of $W$. Therefore we get
Here the last equality comes from $(p-p^{\prime})\in W=\mathrm{span}(\bar{\theta},s)$, thus $\pi_{s^{\perp}}(p-p^{\prime})=\pi_{\bar{\theta}}(p-p^{\prime})$.
Now we focus on the $(d-1)$-dimensional space $s^{\perp}$, and consider the projections $\pi_{s^{\perp}}(a_{1}),\ldots,\pi_{s^{\perp}}(a_{d})$. Since the diameter of $\operatorname{conv}(a_{1},\ldots,a_{d})$ is at most $3$, we have $\max_{i,j\in[d]}\|\pi_{s^{\perp}}(a_{i})-\pi_{s^{\perp}}(a_{j})\|\leq 3$. Because $\theta$ is a unit normal vector of $R$ and $\theta\in s^{\perp}$, we know that $\theta$ is also a unit normal of $\pi_{s^{\perp}}(R)=\pi_{s^{\perp}}(\operatorname{conv}(a_{2},\ldots,a_{d}))$. This gives
Also, since $\operatorname{dist}(F\cap W,\partial R)\geq r$ where $F\cap W=\{p+st:t\in\mathbb{R}\}$ is one-dimensional. After the projection to $s^{\perp}$ we have
$\operatorname{dist}(\pi_{s^{\perp}}(p),\partial\pi_{s^{\perp}}(R))=\operatorname{dist}(F\cap W,\partial R)\geq r.$
Therefore we can use Lemma 37 to get
where the first step comes from $\pi_{s^{\perp}}(p^{\prime})\in\operatorname{affhull}(\pi_{s^{\perp}}(R^{\prime}))$. The lemma then follows from (16). ∎
Lemma 37.
Given $b_{1},\ldots,b_{d}\in\mathbb{R}^{d-1}$ such that $\operatorname{conv}(b_{1},\ldots,b_{d})$ is non-degenerate. Suppose
@@
Then we have $\operatorname{dist}(q,\operatorname{affhull}(b_{1},\ldots,b_{d-1}))\geq r\delta/3$.
Proof.
@@
For simplicity, write $B=\operatorname{conv}(b_{2},\ldots,b_{d})$ and $B^{\prime}=\operatorname{conv}(b_{1},\ldots,b_{d-1})$. Let $q^{\prime}=\pi_{B^{\prime}}(q)$ be the point closest to $q$ on $\operatorname{affhull}(B^{\prime})$, i.e. $\|q-q^{\prime}\|=\operatorname{dist}(q,\operatorname{affhull}(b_{1},\ldots,b_{d-1}))$.
@@
Let $x=(B\cap B^{\prime})\cap\operatorname{affhull}(b_{1},q,q^{\prime})$ be its intersection between the two-dimensional plane $\operatorname{affhull}(b_{1},q,q^{\prime})$ and the $(d-3)$-dimensional ridge $B\cap B^{\prime}$ (which gives a unique point). (See Figure 2 for an illustration). Consider the triangle $\operatorname{conv}(b_{1},q,x)$ and calculate its area in two different ways. On one hand, it has base $\operatorname{conv}(b_{1},x)$ of length $\|b_{1}-x\|\leq 3$ with height $\operatorname{dist}(q,\operatorname{affhull}(b_{1},x))=\|q-q^{\prime}\|$, which gives that the area of the triangle is at most $23\|q-q^{\prime}\|/$ 2. On the other hand, this triangle has base $\operatorname{conv}(x,q)$ of length $\|x-q\|\geq\operatorname{dist}(q,\partial(B))\geq r$ with height $\operatorname{dist}(b_{1},\operatorname{affhull}(x,q))\geq\operatorname{dist}(b_{1},B)\geq\delta$, which gives that the area of the triangle is at least $r\delta/2$.
Combining the above two ways of determining the area of triangle $\operatorname{conv}(b_{1},q,x)$, we have $23\|q-q^{\prime}\|/2\geq r\delta/$ 2. Therefore we have $\operatorname{dist}(q,\operatorname{affhull}(B^{\prime}))=\|q-q^{\prime}\|\geq\frac{r\delta}{3}$ as desired.
[…]
@@
Figure 2: Illustration of Lemma 37 when $d-1=3$. In gray is the intersection between the two-dimensional plane $\operatorname{affhull}(b_{1},q,q^{\prime})$ and $\operatorname{conv}(b_{1},\ldots,b_{d})$. The red triangle is $\operatorname{conv}(b_{1},x,q)$. The bottom face is $B$ and the back face is $B^{\prime}$.
@@ 5.3 Randomized Lower-Bound for $\delta$: Distance between vertices and facets
In this section, we show that the affine hull of a given facet $F$ of the polytope $\operatorname{conv}(a_{1},\ldots,a_{n})$ is $\Omega(\frac{1}{Ld\log n})$-far away to other vertices with good probability, or in other words, the distance $\delta$ in Lemma 36 is at least $\Omega(\frac{1}{Ld\log n})$ with good probability. Our main result of this section is as follows:
Lemma 38 (Randomized lower-bound for $\delta$).
@@
Let $a_{1},\dots,a_{n}\in\mathbb{R}^{d}$ be independent $L$-log-Lipshchitz random vectors. For any $I\in\binom{[n]}{d}$ such that $\Pr[E_{I}]\geq 10\binom{n}{d}^{-1}$, we have
@@
$\Pr[\min_{k\in[n]\backslash I}\operatorname{dist}(\operatorname{affhull}(F_{I}),a_{k})\geq\frac{1}{10e^{3}dL\log n})\mid E_{I}]\geq 0.72.$
@@
To show Lemma 38, we fix any $I\in\binom{[n]}{d}$ of consideration. Without loss of generality, assume $I=[d]$ and write $E=E_{[d]}$. We define the following event $B_{\varepsilon}$ indicating that the distance from $F_{[d]}$ to other vertices is at least $\varepsilon$.
Definition 39 (Separation by the margin of a facet).
@@
Let $\theta\in\mathbb{S}^{d-1},t\in\mathbb{R}$ be as in Definition 18. For any $\varepsilon>0$, let $B_{\varepsilon}^{+}$ denote the event that $\theta^{\top}a_{i}<t-\varepsilon$ for all $i\in[n]\setminus[d]$ and $B_{\varepsilon}^{-}$ denote the event that $\theta^{\top}a_{i}>t+\varepsilon$ for all $i\in[n]\setminus[d]$. We write $B_{\varepsilon}=B_{\varepsilon}^{+}\vee B_{\varepsilon}^{-}$.
@@
Fix any $a_{1},\dots,a_{d}\in\mathbb{R}^{n}$ subject to the non-degeneracy assumptions in 23 and $\operatorname{conv}(a_{1},\dots,a_{d})\cap W\neq\emptyset$. Define $\theta\in\mathbb{S}^{d-1},t>0$ as described in Definition 18, i.e. $\theta^{\top}a_{i}=t$ for each $i\in[d]$. Write $s_{i}=\theta^{\top}a_{i}$ for each $i\in[n]\setminus[d]$. We note that $s_{i}$ is an $L$-log-Lipschitz random variable for all $i\in[n]\backslash[d]$. Moreover, over the remaining randomness in $a_{d+1},\dots,a_{n}$, we have $\Pr[E]=\Pr[B_{0}^{+}]+\Pr[B_{0}^{-}]$ and $\Pr[B_{\varepsilon}]=\Pr[B_{\varepsilon}^{+}]+\Pr[B_{\varepsilon}^{-}]$. We will show that $\Pr[B_{\varepsilon}^{+}]\leq\frac{1}{2\binom{n}{d}}+\frac{5}{4}\Pr[B_{0}^{+}]$, and the appropriate statement will follow for $B_{\varepsilon}^{-}$ analogously. Putting together this will prove the lemma.
If $\Pr[B_{0}^{+}]\leq\frac{1}{2}\binom{n}{d}^{-1}$ then the desired inequality holds directly. Otherwise, fix any $i\in[n]\backslash[d]$ and let $\mu_{i}$ denote the induced probability density function of $s_{i}$. We then have
@@
The first two inequalities above follow from $L$-log-Lipschitzness of $\mu_{i}$. The third follows from the fact that $s_{i}\geq t+1/L$ implies $s_{i}\geq t$. As such we can, for fixed $t,\theta,$ upper-bound the probability over $s_{1},\dots,s_{d}$ that, conditional on $B_{0}^{+}$, there exists a vertex being $\varepsilon$-close to $\operatorname{affhull}(F_{I})$:
@@
To interpret the last equality above, we observe that #$\#\{i\in[n]\setminus[d]:s_{i}\geq t\}=0$ if and only if $B_{0}^{+}$ happens. The upper bound $\Pr[B_{0}^{+}]\leq\frac{5}{4}\Pr[B_{\varepsilon}^{+}]$ then follows from (18) together with 41 and our choice of $\varepsilon$. ∎
Claim 41.
Proof.
$\Pr[X=0]\leq\exp\left(-\frac{\mathbb{E}[X]}{2}\right).$
As such, $\mathbb{E}[X]>2d\log n$ would imply $\Pr[X=0]<n^{-d}$, contradicting the original assumption that $\Pr[X=0]\geq n^{-d}.$ It follows that $\mathbb{E}[X]\leq 2d\log n$. ∎
Now we can prove Lemma 38 using Lemma 40.
Proof of Lemma 38.
Fix any $I\in\binom{[n]}{d}$. By Lemma 40, we have that $\Pr[E_{I}]\leq\binom{n}{d}^{-1}+\frac{5}{4}\cdot\Pr[E_{I}\wedge(\delta\geq\varepsilon)]$ for $\varepsilon=\frac{1}{10e^{3}Ld\log n}$. This gives that

### v3

@@ 5.1 Notations
Since we assume that the constraint matrix rows $a_{1},\dots,a_{n}$ each have a continuous probability density function, $\operatorname{conv}(a_{1},\dots,a_{n})$ and $W$ satisfy the non-degeneracy conditions (see Fact 4) almost surely. In this case, each edge of the polygon $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ is given by $F_{I}\cap W=\operatorname{conv}(a_{i}:i\in I)$ for which $I\in\binom{[n]}{d}$ and $E_{I}$ holds (where $F_{I}$ and $E_{I}$ are defined in Definition 3.1). In addition, each vertex of the polygon $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ is given by the intersection between $W$ and (d2)(d-2)-dimensional ridges of $\operatorname{conv}(a_{1},\ldots,a_{n})$, which are convex hulls of (d1)(d-1) vertices of $\operatorname{conv}(a_{1},\ldots,a_{n})$. We define the following notations for a ridge and its corresponding vertex:
Definition 5.4 (Ridge and vertex event).
@@
Remark 5.5.
Any vertex $v$ of $\operatorname{conv}(a_{1},\dots,a_{n})$ can be written as $v=R_{J}\cap W$ for some $J\subseteq[n]$ for which $A_{J}$ holds. Assuming non-degeneracy, each $J$ for which $A_{J}$ holds satisfies |J|=d1|J|=d-1 and the relation between vertices and index sets $J\in\binom{[n]}{d-1}$ with $A_{J}$ is a one-to-one correspondence.
@@ 5.2 Deterministic Conditions for a Good Edge-to-Vertex Separator
In this subsection, we present a series of sufficient conditions such that an edge on the polygon $\operatorname{conv}(a_{1},\ldots,a_{n})\cap W$ maintains a significant separation from its next vertex in clockwise order. When the vertex set $\{a_{i}\}_{i=1}^{n}$ is fixed, this edge-to-vertex distance can be decomposed into two geometric components:
  • the distance from all other vertices of $Q$ to the supporting hyperplane of the facet containing the edge, and
  • the depth at which the intersection point $p=R\cap W$ lies in the interior of the ridge $R$.
Lemma 5.6 shows that if these two quantities are bounded below by $\delta$ and $r$ respectively, then their product, $r\delta/3$, guarantees a significant edge-to-vertex gap in the projected polygon $Q\cap W$.
Lemma 5.6.
@@
    (Inner radius of $R$) $\operatorname{dist}(\operatorname{affhull}(F\cap W),\partial R)\geq r$.
Then for any $p\in F\cap W$, the outer unit normal vector $\bar{\theta}\in W$ to the edge $F\cap W$ satisfies
$\bar{\theta}^{\top}p-\bar{\theta}^{\top}p^{\prime}\geq\delta r/3,$
where $p^{\prime}\in Q\cap W$ is the next vertex after $F\cap W$ in clockwise order.
The reader who desires a more intuitive illustration of the geometry involved may take a look at Figure 2 from the next lemma, where relevant concepts are depicted as they happen for $d=4$. The simplex is the “next” facet after $F$, and its bottom face $B$ is the ridge of $Q$ that is shared with $F$. The distance from $b_{1}$ to $B$ is large due to the first assumption of Lemma 5.6, and the distance from $q$ to the boundary of $B$ is large due to the second assumption. To map the depicted facet to a three-dimensional space for the purpose of the illustration, it has been projected orthogonally to the subspace perpendicular to $F\cap W$.
We remark that Lemma 5.6 gives a sufficient condition assuming that $\{a_{i}:i\in[n]\}$ is fixed. In later subsections, our goal is to show this sufficient condition actually occurs with high probability even when $\{a_{i}:i\in[n]\}$ are Laplace-Gaussian distributed random variables.
To start, we show a lemma about the distance from a point in the simplex to the affine hull of its neighboring facet.
Lemma 5.7.
Given $b_{1},\ldots,b_{d}\in\mathbb{R}^{d-1}$ such that $\operatorname{conv}(b_{1},\ldots,b_{d})$ is non-degenerate. For any $D>0$, suppose
@@
Then we have $\operatorname{dist}(q,\operatorname{affhull}(b_{1},\ldots,b_{d-1}))\geq r\delta/D$.
Proof 5.8.
@@
For simplicity, write $B=\operatorname{conv}(b_{2},\ldots,b_{d})$ and $B^{\prime}=\operatorname{conv}(b_{1},\ldots,b_{d-1})$. Let $q^{\prime}=\pi_{B^{\prime}}(q)$ be the point closest to $q$ on $\operatorname{affhull}(B^{\prime})$, i.e., $\|q-q^{\prime}\|=\operatorname{dist}(q,\operatorname{affhull}(b_{1},\ldots,b_{d-1}))$.
@@
Let $x=(B\cap B^{\prime})\cap\operatorname{affhull}(b_{1},q,q^{\prime})$ be its intersection between the two-dimensional plane $\operatorname{affhull}(b_{1},q,q^{\prime})$ and the (d3)(d-3)-dimensional ridge $B\cap B^{\prime}$ (which gives a unique point). (See Figure 2 for an illustration).
@@
Figure 2: Illustration of Lemma 5.7 when d1=3d-1=3. In light green is the intersection between the two-dimensional plane $\operatorname{affhull}(b_{1},q,q^{\prime})$ and $\operatorname{conv}(b_{1},\ldots,b_{d})$. The red triangle is $\operatorname{conv}(b_{1},x,q)$. The bottom face is $B$ and the left-facing back face is $B^{\prime}$.
Consider the triangle $\operatorname{conv}(b_{1},q,x)$ and calculate its area in two different ways. On one hand, it has base $\operatorname{conv}(b_{1},x)$ of length $\|b_{1}-x\|\leq D$ with height $\operatorname{dist}(q,\operatorname{affhull}(b_{1},x))=\|q-q^{\prime}\|$, which gives that the area of the triangle is at most $D\|q-q^{\prime}\|/2$. On the other hand, this triangle has base $\operatorname{conv}(x,q)$ of length $\|x-q\|\geq\operatorname{dist}(q,\partial(B))\geq r$ with height $\operatorname{dist}(b_{1},\operatorname{affhull}(x,q))\geq\operatorname{dist}(b_{1},\operatorname{affhull}(B))\geq\delta$, which gives that the area of the triangle is at least $r\delta/2$.
Combining the above two ways of determining the area of triangle $\operatorname{conv}(b_{1},q,x)$, we have $D\|q-q^{\prime}\|/2\geq r\delta/2$. Therefore we have $\operatorname{dist}(q,\operatorname{affhull}(B^{\prime}))=\|q-q^{\prime}\|\geq\frac{r\delta}{R}$ as desired.
What Lemma 5.7 tells us is that, if we have sufficiently strong information about the geometry in base of a simplex (conditions 1 and 2) then we can relate the height of that simplex to the length of a given chord. In our setting, we let a ridge of the polar polytope play the role of the base, and take the apex to be one of the vertices $a_{i}$ not contained in the chosen facet.
To prove our main deterministic geometric result (Lemma 5.6), we squash the configuration by projecting it orthogonally onto the (d1)(d-1)-dimensional subspace $s^{\perp}$ orthogonal to the edge direction. Under this projection, the two neighboring ridges $R$ and $R^{\prime}$ become adjacent facets of a simplex, with the apex corresponding to the first vertex outside the facet $F$. The bounded diameter, height, and inner radius conditions in Lemma 5.6 then align precisely with the assumptions of Lemma 5.7, which allow us to apply it and lift the resulting bound back to the original two-dimensional slice $W$.
Proof 5.9 (Proof of Lemma 5.6).
Figure 3: Illustration of the variables in Lemma 5.6 for $d=3$. The light-green parallelogram is the two-dimensional plane $W$. The red arrows are $\theta$ (outward unit normal of $F$) and $\bar{\theta}=\pi_{W}(\theta)/\|\pi_{W}(\theta)\|$. The red points $p=W\cap R$ and $p^{\prime}=W\cap R^{\prime}$ are two consecutive points on the polygon $Q\cap W$. The red dashed line demonstrates the edge-to-vertex distance $\bar{\theta}^{\top}(p-p^{\prime})$.
Write $R^{\prime}$ for the ridge of $Q$ such that $\{p^{\prime}\}=R^{\prime}\cap W$. Since $p^{\prime}\in Q\cap W$ is adjacent to vertex $p$ and the edge $F\cap W$, by Fact 5 we may relabel the $a_{i}$ such that $R^{\prime}=\operatorname{conv}(a_{1},\dots,a_{d-1})$, $R=\operatorname{conv}(a_{2},\dots,a_{d}),$ and $F=\operatorname{conv}(a_{2},\dots,a_{d+1})$ without loss of generality. Let $\theta\in\mathbb{S}^{d-1}$ denote the outward unit normal to $F$. This normal vector satisfies
Let $s\in\mathbb{S}^{d-1}$ be a unit vector indicating the direction of the (one-dimensional) line $F\cap W$. This vector is unique up to sign. Also, let $\bar{\theta}=\pi_{W}(\theta)/\|\pi_{W}(\theta)\|$ be the outward unit normal to $F\cap W$ in the two-dimensional plane $W$. See Figure 3 for an illustration of the variables for $d=3$. Notice that $\bar{\theta}$ and $s$ form an orthonormal basis of $W$. Therefore we get
Here the last equality comes from $(p-p^{\prime})\in W=\mathrm{span}(\bar{\theta},s)$, so $\pi_{s^{\perp}}(p-p^{\prime})=\pi_{\bar{\theta}}(p-p^{\prime})$.
Now we focus on the (d1)(d-1)-dimensional space $s^{\perp}$, and consider the orthogonal projections $\pi_{s^{\perp}}(a_{1}),\ldots,\pi_{s^{\perp}}(a_{d})$. Since the diameter of $\operatorname{conv}(a_{1},\ldots,a_{d})$ is at most $3$, we have $\max_{i,j\in[d]}\|\pi_{s^{\perp}}(a_{i})-\pi_{s^{\perp}}(a_{j})\|\leq 3$. By definition, $\theta$ is a unit normal of $R$, so since $s\in R$, we have $\theta\in s^{\perp}$. It follows that $\theta$ is also a unit normal of $\pi_{s^{\perp}}(R)=\pi_{s^{\perp}}(\operatorname{conv}(a_{2},\ldots,a_{d}))$. This gives
Also, since $\operatorname{dist}(F\cap W,\partial R)\geq r$ and $F\cap W$ is one-dimensional, after the projection to $s^{\perp}$ we have
$\operatorname{dist}(\pi_{s^{\perp}}(p),\partial\pi_{s^{\perp}}(R))=\operatorname{dist}(\operatorname{affhull}(F\cap W),\partial R)\geq r.$
Therefore we can use Lemma 5.7 to get
where the first step comes from $\pi_{s^{\perp}}(p^{\prime})\in\operatorname{affhull}(\pi_{s^{\perp}}(R^{\prime}))$. The lemma then follows from (16).
@@ 5.3 Randomized Lower-Bound for $\delta$: Distance between vertices and facets
In this section, we show that the affine hull of a given facet $F$ of the polytope $\operatorname{conv}(a_{1},\ldots,a_{n})$ is $\Omega(\frac{1}{Ld\log n})$-far away to other vertices with good probability, or in other words, the distance $\delta$ in Lemma 5.6 is at least $\Omega(\frac{1}{Ld\log n})$ with good probability. Our main result of this section is as follows:
Lemma 5.10 (Randomized lower-bound for $\delta$).
@@
Let $a_{1},\dots,a_{n}\in\mathbb{R}^{d}$ be independent $L$-log-Lipschitz random vectors. For any $I\in\binom{[n]}{d}$ such that $\Pr[E_{I}]\geq 10\binom{n}{d}^{-1}$, we have
$Pr[\min_{k\in[n]\backslash I}\operatorname{dist}(\operatorname{affhull}(F_{I}),a_{k})\geq\frac{1}{10e^{3}dL\log n})\mid E_{I}]\geq 0.72.$
Intuition
The proof of this lemma will span this entire subsection. Let us start with some words on the intuition behind it. Assume $\operatorname{affhull}(F_{I})$ is fixed arbitrarily. Then the quantities $\operatorname{dist}(\operatorname{affhull}(F_{I}),a_{k})$ are determined solely by the points $a_{k}$, $k\in[n]\setminus I$. The points are $L$-log-Lipschitz, which makes each signed distance $\lvert\theta^{\top}a_{k}\rvert=\operatorname{dist}(\operatorname{affhull}(F_{I}),a_{k})\in\mathbb{R}$ into an $L$-log-Lipschitz random variable as well. Any $L$-log-Lipschitz random variable has its probability density function pointwise upper bounded by $L$, and hence the probability that for a given $k\in[n]\setminus I$ we have $\Pr[\theta^{\top}a_{k}\in[-\varepsilon,\varepsilon]]\leq 2\varepsilon L$. A union bound would then give
$Pr[\min_{k\in[n]\backslash I}\operatorname{dist}(\operatorname{affhull}(F_{I}),a_{k})\leq\varepsilon]\leq 2(n-d)\varepsilon L.$
Taking $\varepsilon=\frac{0.28}{2(n-d)L}$ would give us a bound on the probability.
It is clear that this weaker version of Lemma 5.10 would be relatively easy to prove. However, it has a linear dependence on $n$ and thus it would add a factor of $\sqrt{n}$ to our shadow bound. This is undesirable. To obtain the stronger conclusion, we consider the expected number of points at close distance to $\operatorname{affhull}(F_{I})$.
@@
$Pr\Big[\min_{k\in[n]\backslash I}\operatorname{dist}(\operatorname{affhull}(F_{I}),a_{k})\leq\varepsilon~\Big|~E_{I}\Big]\leq\mathbb{E}\Big[\lvert\{i\in[k]\setminus I:\operatorname{dist}(\operatorname{affhull}(F_{I}),a_{k})\leq\varepsilon\}\rvert~\Big|~E_{I}\Big].$ (17)
If $\varepsilon\leq 1/L$ then we can use $L$-log-Lipschitzness to derive a lower bound on the number of points $a_{k}$ lying above (or below) the affine subspace $\operatorname{affhull}(F_{I})$. What we prove is that if (17) is large then (without conditioning on $E_{I}$) both the expected number of points above $\operatorname{affhull}(F_{I})$ and the expected number of points below $\operatorname{affhull}(F_{I})$ are at least $\Omega(\frac{1}{\varepsilon L})$ times as large as (17).
However, recall that the event $E_{I}$ requires all points to lie on the same side of $\operatorname{affhull}(F_{I})$. If there is simultaneously a point above $\operatorname{affhull}(F_{I})$ and a point below $\operatorname{affhull}(F_{I})$ then $E_{I}$ does not hold. Using the Chernoff bound we can show, conditional on some $\operatorname{affhull}(F_{I})=H$, that if we have $\Pr[E_{I}\mid\operatorname{affhull}(F_{I})=H]\geq n^{-d}$ then at least one of the expected number of points above $\operatorname{affhull}(F_{I})$ or the expected number of points below $\operatorname{affhull}(F_{I})$ must be at most $2d\log n$. If one of these is bounded from above, then (17) must be bounded from above. Taking proper care to observe that those affine subspaces $H$ for which $\Pr[E_{I}\mid\operatorname{affhull}(F_{I})=H]\geq n^{-d}$ together account for most of the probability mass, this will yield the desired result.
Proving the lemma
@@
To show Lemma 5.10, we fix any $I\in\binom{[n]}{d}$ of consideration. Without loss of generality, assume $I=[d]$ and write $E=E_{[d]}$. We define the following event $B_{\varepsilon}$ indicating that the distance from $\operatorname{affhull}(F_{[d]})$ to other vertices is at least $\varepsilon$.
Definition 5.11 (Separation by the margin of a facet).
@@
Let $\theta\in\mathbb{S}^{d-1},t\in\mathbb{R}$ be as in Definition 2.12. For any $\varepsilon>0$, let $B_{\varepsilon}^{+}$ denote the event that $\theta^{\top}a_{i}<t-\varepsilon$ for all $i\in[n]\setminus[d]$ and $B_{\varepsilon}^{-}$ denote the event that $\theta^{\top}a_{i}>t+\varepsilon$ for all $i\in[n]\setminus[d]$. We write $B_{\varepsilon}=B_{\varepsilon}^{+}\vee B_{\varepsilon}^{-}$.
@@
Fix any $a_{1},\dots,a_{d}\in\mathbb{R}^{n}$ subject to $\operatorname{conv}(a_{1},\dots,a_{d})\cap W\neq\emptyset$ and the non-degeneracy assumptions in Fact 4. Define $\theta\in\mathbb{S}^{d-1},t>0$ as described in Definition 2.12, i.e., $\theta^{\top}a_{i}=t$ for each $i\in[d]$. Write $s_{i}=\theta^{\top}a_{i}$ for each $i\in[n]\setminus[d]$. We note that $s_{i}$ is an $L$-log-Lipschitz random variable for all $i\in[n]\backslash[d]$. Moreover, over the remaining randomness in $a_{d+1},\dots,a_{n}$, we have $\Pr[E]=\Pr[B_{0}^{+}]+\Pr[B_{0}^{-}]$ and $\Pr[B_{\varepsilon}]=\Pr[B_{\varepsilon}^{+}]+\Pr[B_{\varepsilon}^{-}]$. We will show that
$\Pr[B_{0}^{+}]\leq\frac{1}{2\binom{n}{d}}+\frac{5}{4}\Pr[B_{\varepsilon}^{+}]$ (18)
and the appropriate statement will follow for $B_{\varepsilon}^{-}$ analogously. Putting together this will prove the lemma. If $\Pr[B_{0}^{+}]\leq\frac{1}{2}\binom{n}{d}^{-1}$ then the desired inequality holds directly.
In order to prove (18) we require the following claim:
Claim 6.
Proof 5.14.
$\Pr[X=0]\leq\exp\left(-\frac{\mathbb{E}[X]}{2}\right).$
As such, $\mathbb{E}[X]>2d\log n$ would imply $\Pr[X=0]<n^{-d}$, contradicting the original assumption that $\Pr[X=0]\geq n^{-d}.$ It follows that $\mathbb{E}[X]\leq 2d\log n$.
Thus, in what remains, we may suppose that $Pr[B_{0}^{+}]>1/2\binom{n}{d}^{-1}$. Fix any $i\in[n]\backslash[d]$ and let $\mu_{i}$ denote the induced probability density function of $s_{i}$. We then have a sequence of inequalities as found below. The first two inequalities above follow from $L$-log-Lipschitzness of $\mu_{i}$. For the first inequality in particular, note that for any $s\in[-\frac{1}{L},0]$, we have $\lvert s-\varepsilon Ls\rvert\leq 1/L$ from $\varepsilon\leq 1/L$. This then gives $\frac{\mu_{i}(t+\varepsilon Ls)}{\mu_{i}(t+s)}\leq\exp(L\cdot(1-\varepsilon L)s)\leq 1$.
@@
The third inequality, on the final line, follows from the fact that $s_{i}\geq t+1/L$ implies $s_{i}\geq t$, and hence $\Pr[s_{i}\geq t\mid s_{i}>t+1/L]=1$. As such we can, for fixed $t,\theta,$ upper-bound the probability over $s_{1},\dots,s_{d}$ that, conditional on $B_{0}^{+}$, there exists a vertex being $\varepsilon$-close to $\operatorname{affhull}(F_{I})$:
@@
To interpret the last equality above, we observe that $\{i\in[n]\setminus[d]:s_{i}\geq t\}=0$ if and only if $B_{0}^{+}$ happens. Then by applying (20) to Claim 6 (note that we are using the assumption $Pr[B_{0}^{+}]>1/2\binom{n}{d}^{-1}$) with our choice of $\varepsilon$ we conclude that
$\Pr[B_{0}^{+}]\leq\frac{5}{4}\Pr[B_{\varepsilon}^{+}]\leq\frac{1}{2}\binom{n}{d}^{-1}+\frac{5}{4}\Pr[B_{\varepsilon}^{+}].$
Now we can prove Lemma 5.10 using Lemma 5.12.
Proof 5.15 (Proof of Lemma 5.10).
Fix any $I\in\binom{[n]}{d}$. By Lemma 5.12, we have that $\Pr[E_{I}]\leq\binom{n}{d}^{-1}+\frac{5}{4}\cdot\Pr[E_{I}\wedge(\delta\geq\varepsilon)]$ for $\varepsilon=\frac{1}{10e^{3}Ld\log n}$. This gives that